If x^2 -7x = 0, the value of x is
1
Expert's answer
2020-08-17T19:16:05-0400

We can factor "x^2-7x=x(x-7)=0", one of the factors "x" and "x-7" must be equal , hence "x=0" or "x-7=0", therefore, "x=0" or "x=7."

Answer: "x_{1}=0,\ x_2=7."